# Controlling Whisper: Universal Acoustic Adversarial Attacks to Control Speech Foundation Models

### v1

@@ 1 Introduction
To highlight the susceptibility of multi-task ASR models to such control-attacks, this work proposes a practical method to subvert the ‘transcription’ setting of the Whisper model, and instead encouraging Whisper to execute speech ‘translation’. Given that an adversary typically does not have access to the textual decoder prompt, we adopt a threat model where the attack is restricted to the acoustic space. Our findings reveal that a short (less than 3 seconds) universal adversarial acoustic segment can be prepended to any speech signal and override Whisper’s ‘transcribe’ setting and execute the ‘translate’ function instead (depicted in Figure 1). Experimental evaluations across four languages indicate that this universal acoustic attack segment can consistently manipulate Whisper’s behavior for nearly all test samples. Further examination of more constrained attack scenarios reveals an interesting bi-modal pattern: the attacks are either highly successful in overriding the command or entirely ineffective for specific samples.
@@ 2.2 Audio Adversarial Attacks
However, with the emergence of the multi-task speech enabled foundation models, in this work, we demonstrate that there is the threat of a new kind of adversarial attack: model-control attack. With the flexibility to perform multiple tasks, an adversary can override the operational setting of a deployed ASR model and force it to perform a different target task - i.e., take control of the model. With Whisper, we illustrate that despite being deployed in ‘transcribe’ mode, a simple universal adversarial attack can always force Whisper to operate in ‘translate’ model.
@@ 5.1 Experimental Setup
Attack Configuration. As described in Section 4, we learn a universal adversarial acoustic segment that can be prepended to any input speech signal to control Whisper to perform speech translation, despite Whisper being deployed in the speech transcription setting. We present experimental results for Whisper medium (769M parameters). Smaller models are not considered as it is found that due to their lower performance in speech translation it limits the potential of the model-control attack (the model-control attack assumes the model has the ability to perform the target task). Following the imperceptibility definitions given in Section 4.2, we learn universal adversarial acoustic segments of three different strengths: 1) a weak attack (attack-w) with the harshest constraints of a maximum amplitude and an audio length of 10,240 frames, equivalent to 0.64 seconds for audio sampled at 16kHz; 2) a mid strength attack (attack-m) with and length 0.64 seconds; and 3) a strong attack (attack-s) with and length 2.56 seconds.
@@ 5.2 Attack Results and Analysis
From 2, in general the greater the probability of English in the generated translation, the higher the quality of the translation as measured by BLEU. However, the curves also reveal a more interesting behaviour of the attacks: the success recall curves (2a), have an observable discontinuity in the BLEU scores when 43%, 62% and 98% of the samples are ‘successfully’ attacked for the weak, mid and strong attacks respectively, and conversely in the fail recall curves (2b), there is a discontinuity when approximately 57%, 38% and 2% samples ‘fail’ to be attacked. These percentages almost perfectly align with the average probability of English given by each attack in Table 1. This suggests that the attacks result in a highly bi-modal response, where the attack on a specific input audio sample is either perfectly successful in causing Whisper to generate English text or perfectly unsuccessful, with near 0% probability of English. This bi-modal split is further verified in the distribution of P(en) given for each attack in Figure 3. When Whisper is running in transcribe (tc) mode, it is entirely non-English (left-most bar), whilst when Whisper is running in translate (tl) mode, it generates entirely English text (right-most bar). However for all of the attacks in Figure 3, we observe the above-mentioned bi-modal pattern, where the model-control attacks do not result in a gradual distributional shift from the French transcriptions to English translations, but instead display binary success. The stronger the attack the greater the proportion of the samples that are ‘successfully’ attacked (greater fraction of samples on the right side) resulting in almost complete English text. On the whole, the universal model-control attack, when successful, causes Whisper to behave as though as if it is in translation mode, but when the attack is unsuccessful it results in Whisper continuing to behave as set in its default transcription mode - there are no partial modes. It is the strictness around the imperceptibility constraints (strength of the attack) that dictates the fraction of samples the universal model-control attack is successful upon.
@@ 5.3 Language Ablation
Next, to verify that the attacks on de/ru/ko display the same bi-modal behaviour as the attack on fr-en, we consider the BLEU-recall curves for each language pair in Figure 4. Most easily observable in Figure 4b the discontinuity in BLEU performance occurs as before with fr-en at the percentage of samples corresponding to approximately the average probability of not English (in Table 2) - between 2% and 5% for the different languages. This verifies that for all languages the same trend holds: the model-control attack is highly binary in either successfully overriding Whisper to translate the source sample to complete English, or failing completely, such that the source audio is not translated at all.
@@ 6 Conclusion
This work reveals a significant vulnerability in multi-tasking speech-enabled foundation models, specifically through model-control adversarial attacks. Our study demonstrates that it is possible to manipulate the task setting of such models, as evidenced by our experiments with OpenAI’s Whisper model. By appending a short universal adversarial acoustic segment to any input speech signal, we were able to override the model’s prompt setting and force it to perform speech translation instead of its default speech transcription. An intriguing aspect is the binary nature of the attack’s success. The universal adversarial acoustic segment either successfully manipulates Whisper to operate in its translation mode, or it fails entirely, resulting in Whisper behaving in its default transcription mode. This bi-modal pattern indicates that the attack does not create intermediate operational states, but rather a clear switch between transcription and translation. The strictness around the imperceptibility constraints dictates the fraction of samples the universal model-control attack switches. The risk of model-control attacks highlights the need for increased security measures in the deployment of flexible ASR systems. As these models are designed to perform an increasing number of tasks within a single framework, they become more susceptible to such adversarial manipulations. We encourage future work on developing robust defenses against model-control adversarial attacks.

### v2

# Controlling Whisper: Universal Acoustic Adversarial Attacks to Control Multi-task Automatic Speech Recognition Models

@@ 1 Introduction
To highlight the susceptibility of multi-task ASR models to such control-attacks, this work proposes a practical method to subvert the ‘transcription’ setting of the Whisper model, and instead encouraging Whisper to execute speech ‘translation’. Given that an adversary typically does not have access to the textual decoder prompt, we adopt a threat model where the attack is restricted to the acoustic space. Our findings reveal that a short (5 seconds) universal adversarial acoustic segment can be prepended to any speech signal and override Whisper’s ‘transcribe’ setting and execute the ‘translate’ function instead (depicted in Figure 1). Experimental evaluations across four languages indicate that this universal acoustic attack segment can consistently manipulate Whisper’s behavior for nearly all test samples. Further examination of more constrained attack scenarios reveals an interesting bi-modal pattern: the attacks are either highly successful in overriding the command or entirely ineffective for specific samples.
@@ 2.2 Audio Adversarial Attacks
However, with the emergence of the multi-task speech enabled foundation models, in this work, we demonstrate that there is the threat of a new kind of adversarial attack: model-control attack. With the flexibility to perform multiple tasks, an adversary can override the operational setting of a deployed ASR model and force it to perform a different target task - i.e., take control of the model. With Whisper, we illustrate that despite being deployed in ‘transcribe’ mode, a simple universal adversarial attack can always force Whisper to operate in ‘translate’ mode.
@@ 5.1 Experimental Setup
Attack Configuration. As described in Section 4, we learn a universal adversarial acoustic segment that can be prepended to any input speech signal to control Whisper to perform speech translation, despite Whisper being deployed in the speech transcription setting. We present experimental results for Whisper medium (769M parameters). Smaller models are not considered as it is found that due to their lower performance in speech translation it limits the potential of the model-control attack (the model-control attack assumes the model has the ability to perform the target task). Following the imperceptibility definitions given in Section 4.2, we learn universal adversarial acoustic segments of three different strengths: 1) a weak attack (attack-w) with the harshest constraints of a maximum amplitude and an audio length of 10,240 frames, equivalent to 0.64 seconds for audio sampled at 16kHz; 2) a mid strength attack (attack-m) with and length 0.64 seconds; and 3) a strong attack (attack-s) with and length 5.12 seconds.
@@ 5.2 Attack Results and Analysis
From 2, in general the greater the probability of English in the generated translation, the higher the quality of the translation as measured by BLEU. However, the curves also reveal a more interesting behaviour of the attacks: the success recall curves (2a), have an observable discontinuity in the BLEU scores when 43%, 62% and 98% of the samples are ‘successfully’ attacked for the weak, mid and strong attacks respectively, and conversely in the fail recall curves (2b), there is a discontinuity when approximately 57%, 38% and 2% samples ‘fail’ to be attacked. These percentages almost perfectly align with the average probability of English given by each attack in Table 1. This suggests that the attacks result in a highly bi-modal response, where the attack on a specific input audio sample is either perfectly successful in causing Whisper to generate English text or perfectly unsuccessful, with near 0% probability of English. This bi-modal split is further verified in the distribution of P(en) for each attack in Figure 3. With Whisper in transcribe (tc) mode, it is entirely non-English (left-most bar), whilst when Whisper in translate (tl) mode, it generates entirely English text (right-most bar). However for all of the attacks in Figure 3, we observe the above-mentioned bi-modal pattern, where the model-control attacks do not result in a gradual distributional shift from the French transcriptions to English translations, but instead display binary success. The stronger the attack the greater the proportion of the samples that are ‘successfully’ attacked (greater fraction of samples on the right side) resulting in almost complete English text. On the whole, the universal model-control attack, when successful, causes Whisper to behave as though as if it is in translation mode, but when the attack is unsuccessful it results in Whisper continuing to behave as set in its default transcription mode - there are no partial modes. It is the strictness around the imperceptibility constraints (strength of the attack) that dictates the fraction of samples the universal model-control attack is successful upon.
@@ 5.3 Language Ablation
Next, to verify that the attacks on de/ru/ko display the same bi-modal behaviour as the attack on fr-en, we consider the BLEU-recall curves for each language pair in Figure 4. Most easily observable in Figure 4b the discontinuity in BLEU performance occurs as before with fr-en at the percentage of samples corresponding to approximately the average probability of not English (in Table 2) - between 2% and 5% for the different languages. This verifies that for all languages the same trend holds: the model-control attack is highly binary.
@@ 6 Conclusion
This work reveals a vulnerability in multi-tasking speech-enabled foundation models, specifically through model-control adversarial attacks. We demonstrate that it is possible to manipulate the task setting of such multi-tasking speech-enabled models. By appending a short universal adversarial acoustic segment to any input speech signal, we were able to override the model’s prompt setting and force it to perform speech translation instead of its default speech transcription. An intriguing aspect is the binary nature of the attack’s success. The universal adversarial acoustic segment either successfully manipulates Whisper to operate in its translation mode, or it fails entirely, resulting in Whisper behaving in its default transcription mode. This bi-modal pattern indicates that the attack does not create intermediate operational states, but rather a clear switch between transcription and translation. The strictness around the imperceptibility constraints dictates the fraction of samples the universal model-control attack switches. The risk of model-control attacks highlights the need for increased security measures in the deployment of flexible ASR systems. As these models are designed to perform an increasing number of tasks within a single framework, they become more susceptible to such adversarial manipulations. We encourage future work on developing robust defenses against model-control adversarial attacks. 222Acknowledgement: We thank Cambridge University Press & Assessment for their support in funding this work.